Find the error in this proof that 2=1.
x=y
X^2=xy
X^2-y^2=xy-y^2
(x+y)(x-y)=y(x-y)
((x+y)(x-y)/(x-y))=(y(x-y)/(x-y))

X+y=y
Y+y=y
2y=y
(2y/y)=(y/y)
2=1

Answer by Mathpassionate(25)



The error is here:
((x+y)(x-y)/(x-y))=(y(x-y)/(x-y))
Because we know that x=y
So x-y = 0
And we cannot divide by zero.